# Experimental Ground-State Energy of a 125-Site Flat Kagome Antiferromagnet via Hamiltonian Engineering on Quantum Computer

### v2

@@ VI Benchmarking Against the Thermodynamic Ground State Energy
Moreover, this analysis reinforces the potential for scaling our method to larger kagome lattices, where the proportion of boundary sites diminishes and finite-size effects are further suppressed. In such bulk-dominated systems, the accuracy of VQE-derived ground-state energy estimates is expected to improve, strengthening the case for quantum simulation of frustrated spin systems using near-term quantum devices.
@@ VII Conclusion
The methods and experimental results presented in this study demonstrate a scalable and physically grounded approach for estimating the ground-state energy of the spin-$\frac{1}{2}$ kagome antiferromagnetic Heisenberg (KAFH) model on large general Kagome lattices. By employing Hamiltonian engineering, specifically, increasing the exchange coupling strength on selected bonds within defect triangles, we circumvent several limitations commonly encountered in prior studies. These include constraints on system size, dimensionality, circuit depth, and the expressibility of the variational quantum eigensolver (VQE) ansatz. Importantly, our approach alleviates the need for preparing quantum states with complex, long-range entanglement, which remains a significant challenge for current hardware architectures.
Our use of enhanced exchange coupling is not only a computational convenience, but also deliberate strategy that leverages core features of quantum spin liquids, such as resonating valence bond (RVB) physics, loop flips, and the quantum superposition of multiple dimer coverings. By initializing the VQE with a physically motivated valence bond crystal (VBC) configuration, we guide the optimizer toward low-energy regions of the Hilbert space while maintaining hardware efficiency. This hybrid strategy—melding insights from many-body physics with the practical constraints of noisy intermediate-scale quantum (NISQ) devices—enables an accurate and resource-conscious estimation of the ground-state energy on systems as large as 125 sites.
@@ Acknowledgements.
We acknowledge the use of IBM Quantum services, including IBM Quantum Credits for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. Our experiments used following IBM quantum processors: ibm_aachen, ibm_kingston, ibm_torino, ibm_fez and ibm_marrakesh.
@@ Appendix A Supplementary material
Figure 6: (Color online) (a) VQE ansatz circuit with qubit indices mapped to the corresponding lattice sites on the Kagome lattice shown in Figure 2. (b) Physical qubit layout of the ibm_falcon processor, showing how circuit qubits are assigned to hardware qubits. The complete mapping can be followed by tracing the red outline from qubit 0 through 1, 2, $\ldots$, up to 100, 101, and 102.
@@
Figure 8: (Color online) Subregions of the Kagome lattice used in local VQE, with corresponding ground-state energy components listed in Table 2. Numerical labels on the lattice sites indicate qubit indices from the ansatz circuit shown in Figure 6(a). The qubit mapping defines a specific dimer covering (thick black line segments), and highlights the defect triangle (in red). Table 2 also lists the bonds selected for Hamiltonian engineering within each defect triangle, along with their enhanced coupling values $J^{\prime}$.

### v3

@@ VI.1 Open Boundary Correction
This analysis reinforces the potential for scaling our method to larger kagome lattices, where the proportion of boundary sites diminishes and finite-size effects are further suppressed. In such bulk-dominated systems, the accuracy of VQE-derived ground-state energy estimates is expected to improve, strengthening the case for quantum simulation of frustrated spin systems using near-term quantum devices.
@@ VII Conclusion
The methods and experimental results presented in this study demonstrate a scalable and physically grounded approach for estimating the ground-state energy of the spin-$\frac{1}{2}$ kagome antiferromagnetic Heisenberg model on large general Kagome lattices. By employing Hamiltonian engineering, specifically, increasing the exchange coupling strength on selected bonds within defect triangles, we circumvent several limitations commonly encountered in prior studies. These include constraints on system size, dimensionality, circuit depth, and the expressibility of the variational quantum eigensolver ansatz. Importantly, our approach alleviates the need for preparing quantum states with complex, long-range entanglement, which remains a significant challenge for current hardware architectures.
Our use of enhanced exchange coupling is not only a computational convenience, but also deliberate strategy that leverages core features of quantum spin liquids, such as resonating valence bond (RVB) physics, loop flips, and the quantum superposition of multiple dimer coverings. By initializing the VQE with a physically motivated valence bond crystal (VBC) configuration, we guide the optimizer toward low-energy regions of the Hilbert space while maintaining hardware efficiency. This hybrid strategy, melding insights from many-body physics with the practical constraints of NISQ devices, enables an accurate and resource-conscious estimation of the ground-state energy on systems as large as 125 sites.
@@
We provide our implementation as open-source to support transparency and quantum computing community engagement at https://github.com/ahsan-quantum/Kagome-Lattice-VQE.git.
@@ Acknowledgements.
We acknowledge the use of IBM Quantum services, including IBM Quantum Credits for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. Our experiments used following IBM quantum processors: ibm_aachen, ibm_kingston, ibm_torino, ibm_fez and ibm_marrakesh. We also acknowledge the support of Higher Education Commission Pakistan under the project National Center of Quantum Computing.
@@ Appendix A Supplementary material
Figure 6: (Color online) (a) VQE ansatz circuit with qubit indices mapped to the corresponding lattice sites on the Kagome lattice shown in Figure 2. (b) Physical qubit layout of the IBM Heron r1 processor (ibm_kingston), showing how circuit qubits are assigned to hardware qubits. The complete mapping can be followed by tracing the red outline from qubit 0 through 1, 2, $\ldots$, up to 100, 101, and 102.
@@
Figure 8: (Color online) Subregions of the Kagome lattice of Figure-2 used in local VQE, with corresponding ground-state energy components listed in Table 2. Numerical labels on the lattice sites indicate qubit indices from the ansatz circuit shown in Figure 6(a). The qubit mapping defines a specific dimer covering (thick black line segments), and highlights the defect triangle (in red). Table 2 also lists the bonds selected for Hamiltonian engineering within each defect triangle, along with their enhanced coupling values $J^{\prime}$.